Draw the net electric force on a +1-C charge placed at the center of:

A line segment of length 1.0 m with a +1-C charge and a –1-C charge at its endpoints;


1
Expert's answer
2021-06-13T18:05:23-0400

Force on center charge due to 1C toward negative charge, is,

"F=\frac{1}{4\pi\epsilon_0}\frac{q_1\times q_2}{r^2}=\frac{9 \times 10^9 \times 1 \times 1}{0.25}=36 \times 10^9N"


Force due to -1C will be attractive, towards -1C charge, is

"F'=\frac{1}{4\pi\epsilon_0}\frac{q_3\times q_2}{r^2}=\frac{9 \times 10^9 \times 1 \times 1}{0.25}=36 \times 10^9N"


This force will act towards the negative charge.




So, total force on center charge will be,

"F_T=F+F'=(36 \times 10^9) +( 36 \times 10^9 ) = 72 \times 10^9N"